# EAMA : Entity-Aware Multimodal Alignment Based Approach for News Image Captioning

### v4

@@ Abstract.
News image captioning requires model to generate an informative caption rich in entities, with the news image and the associated news article. Though Multimodal Large Language Models (MLLMs) have demonstrated remarkable capabilities in addressing various vision-language tasks, our research finds that current MLLMs still bear limitations in handling entity information on news image captioning tasks. Besides, while MLLMs have the capability to process long inputs, generating high-quality news image captions still requires a trade-off between sufficiency and conciseness of textual input information. To explore the potential of MLLMs and address problems we discovered, we propose EAMA: an Entity-Aware Multimodal Alignment based approach for News Image Captioning. Our approach first aligns the MLLM through Balance Training Strategy with two extra alignment tasks: Entity-Aware Sentence Selection task and Entity Selection task, together with News Image Captioning task, to enhance its capability in handling multimodal entity information. The aligned MLLM will utilizes the additional entity-related information it explicitly extract to supplement its textual input while generating news image captions. Our approach achieves better results than all previous models in CIDEr score on GoodNews dataset (72.33 -¿ 88.39) and NYTimes800k dataset (70.83 -¿ 85.61).
Multimodal Large Language Model, News Image Captioning, Multimodal Alignment, Multimodal Entity Information
copyright: noneccs: Computing methodologies Natural language processingccs: Computing methodologies Computer visionccs: Computing methodologies Computer vision tasks
1. Introduction
@@
News image captioning, a variant of image captioning task that aims to generate the image caption which contains more specific information, has developed a lot as a standard multimodal task in recent years(Zhang et al., 2022; Rajakumar Kalarani et al., 2023; Qu et al., 2023). In the news image captioning task, models are required to generate an informative caption, rich in entities such as the names of people or news events, based on the provided news image and associated news article, as illustrated in the example shown in Figure 1. According to (Tran et al., 2020), over 96% news image captions contain entity information and approximately 26% words of news image captions are entities in common news image captioning datasets. Compared with general image captioning task, news image captioning task requires model to not only understand the overall information of images, but also apply the knowledge of model to extract related specific information from associated news articles.
@@
Figure 1. An example of news image caption task and response of InstructBLIP in different settings. Origin denotes Zero_Shot, OSFT denotes finetuning InstructBLIP in official settings.
@@
Recently, multimodal large language models are developing rapidly and have shown promising capability to address traditional image captioning task. Therefore as a variant of image captioning task, it is natural to consider MLLMs as a promising technique for the news image captioning task for two reasons. What’s more, news image captioning task requires extracting related information from a long news article, which is exactly a task that can be handled by Large Language models because they contains extensive knowledge and exhibits powerful text processing ability. Since MLLMs mostly bear a LLM part, we believe MLLMs is potentially good at news image captioning task.
@@
However, through our experiments, we find that current MLLMs such as InstructBLIP still face some challenges on news image captioning task: (1) Current MLLMs are not good at dealing with multimodal entity information in zero-shot setting. The ability of generating entities remains insufficient after simply finetuned on news image captioning dataset. Figure 1 shows a specific example. (2) While MLLMs are capable of processing long textual information and can take almost entire news articles as textual input, the generation of news image captions still requires a trade-off between sufficiency and conciseness of textual input information.
@@
To address the problems we discovered, we propose EAMA, an Entity-Aware Multimodal Alignment based approach for MLLMs. Specifically, we first design two entity-aware alignment tasks, which are Entity-Aware Sentence Selection task and Entity Selection task. Then we align the MLLM with original news image captioning task and two alignment tasks applying our Balance Training Strategy to promote the MLLM’s ability of handling multimodal entity information. Finally, we utilize the aligned MLLM to extract for some relevant information from the entire news article to supplement the textual input context. Our method not only obtains the best result on all automatic metrics but also achieves the best performance on entity generation.
Figure 2. An overview of our approach: EAMA. The upper part represents Entity-Aware Multimodal Alignment with Entity-Aware Sentence Selection, Entity Selection and News Image Captioning task through Balance Training Strategy. The bottom part represents the Self-Supplemented Generation for the news image caption given the news article and the news image.
Our contributions are:
  • We find that current MLLMs are poor at dealing with mutlimodal entity information under zero-shot setting. The ability remains insufficient after simply finetuned on news image captioning dataset. Besides, generating high-quality news image captions with the MLLM still requires a trade-off between sufficiency and conciseness of textual input information.
  • We design two alignment tasks and align the MLLM with our Balance Training Strategy to promote the MLLM’s ability for handling entity multimodal information. Besides, the aligned MLLM will utilize the additional entity-related information they explicitly extract to supplement their textual input while generating news image captions.
  • Our Entity-Aware Multimodal Alignment based approach: EAMA, outperforms previous methods on both GoodNews and NYTimes800k Datasets.
@@
Our code and data will be released to the community to facilitate future research.
2. Related Works
2.1. News Image Captioning Task
News image captioning(Biten et al., 2019; Tran et al., 2020; Liu et al., 2021) is a variant task of the general image captioning task which takes news images and associated news articles as input to generate informative captions of news images. In early researches, models((Lu et al., 2018), (Biten et al., 2019)(Avg+CtxIns)) often generate templates which do not contain specific entities, and then extract entities from news articles. However, template-based methods often constrain the representation of descriptions. Newly proposed end-to-end methods(Tran et al., 2020; Yang et al., 2021; Zhang et al., 2022; Rajakumar Kalarani et al., 2023; Qu et al., 2023) generate the final caption in one single step and achieve higher performance. Tell(Tran et al., 2020) applies RoBERTa and ResNet-152 to encode textual and visual information and uses transformer decoder to generate captions. JoGANIC (Yang et al., 2021) is a transformer based model with component template guidance following the journalistic guidelines. NewsMEP (Zhang et al., 2022) applies CLIP (Radford et al., 2021) to translate the image with open-domain multimodal knowledge and BART to encode textual and vision information simultaneously. Rajakumar Kalarani et al. (2023) pretrains OFA model on three vision-language tasks which requires extra data, and then finetunes OFA on news image captioning datasets. Qu et al. (2023) also utilizes BART architecture and design distinct modules for different vision inputs.
However, few researches with MLLMs have been done in news image captioning task while MLLMs are powerful in various multimodal tasks. As for alignment, while Qu et al. (2023) leverages a weakly supervised face-name alignment method to refine face representations, it only focuses on people’s names and ignores other types of entities. And its alignment method requires modifying the BART architecture and necessitates that the image contains a facial region as they need to first identify the face region. Our approach is the first to align the MLLM with newly designed alignment tasks focusing on general types of entities. Our approach utilizes all data constructed from the corresponding news image caption dataset without additional data or modifications to the structure of MLLMs.
Since news articles are too long for transformer-based model in previous research, many methods have been designed to address the problem. JoGANIC(Yang et al., 2021) proposes a Multi-Span Text Reading (MSTR) method to read more information from articles. ICECAP(Hu et al., 2020) trains an extra VSE++ model to retrieve relevant sentences where entities are replaced with their corresponding categories and then generate the template caption and predict the position of entity tokens in extracted relevant sentences. Zhou et al. (2022a) utilizes CLIP model trained on news image caption data and OpenNRE to obtain the key local context to generate the news image captions.
However, MSTR requires much longer input tokens and all retrieval-based methods require an extra cross-modal retrieval model as well as an extra training cost while the capability of the model to generate news image captions itself has not been improved. Our approach aligns the MLLM to promote its ability on handling multimodal entity information. The aligned MLLM can then be used to explicitly extract entity-related sentences to supplement the textual input context without extra models or training costs .
2.2. Multimodal Large Language Models and Alignment
@@
Multimodal large language models have developed rapidly in recent years. BLIP-2(Li et al., 2023a) applys Q-Former architecture to transform image input into LLMs input tokens. LLaVA(Liu et al., 2023b) and LLaVA-v1.5(Liu et al., 2023a) utilize linear layers or perceptrons to map the vision features into the inputs of LLMs. Through instruction tuning on BLIP2, InstructBLIP(Dai et al., 2023) gains the ability to follow the instruction on different tasks. There are many other MLLMs such as mPLUG-Owl(Ye et al., 2023), Otter(Li et al., 2023b) and Qwen-VL (Bai et al., 2023). Among all MLLMs, GPT-4V(OpenAI et al., 2023) is the most powerful one now. We select some of these MLLMs on our research.
@@
There are numerous studies on multimodal alignment tasks recent years. The most traditional task is image-text matching in CLIP(Radford et al., 2021). There are also fine-grained alignment tasks such as Region-Tag and Region-Noun alignment tasks in Zhou et al. (2022b). MLLMs such as (Li et al., 2023a; Ye et al., 2023) often align textual features and visual features on multimodal tasks with language modeling loss.
3. Approach
3.1. Overview
@@
MLLMs (Multimodal Large Language Models) prove highly effective across diverse multimodal tasks. Our methodology is crafted around MLLMs, informing both our approach design and experimental framework. Figure 2 offers a concise overview, delineating the alignment training and caption generation stages. Typically, MLLM architectures comprise an image encoder, a Vision-Language (V-L) Connector, and a large language model.
@@
During our alignment training stage, we commence by simultaneously training the MLLM on three tasks: Entity-Aware Sentence Selection, Entity Selection, and News Image Captioning. This approach employs a balanced training strategy aimed at enhancing the MLLM’s capacity to handle multimodal entity information effectively.
@@
We employ a language modeling loss as the optimization objective across all tasks. Typically, within a given dataset $D$, each data instance comprises three components: the input image $I$, the input context $T$, and the target output text $O=(o_{1},...,o_{n})$, with $o_{i}$ denoting a token. Thus, the general multimodal language modeling loss can be formulated as:
(1) $\centering\mathcal{L}_{\mathrm{LM}}([I;T;O])=-\dfrac{1}{n}\sum_{i=1}^{n}\log\mathbf{P}(o_{i}|o_{1},...,o_{i-1};I;T)\@add@centering$
@@
For example, in News Image Captioning Task, the input image $I$ is the news image, the input context $T$ is the news article, while the target output text $O$ is the given caption. Thus for a News Image Captioning dataset $D$, the corresponding News Image Captioning loss can be denoted as:
(2) $\centering\mathcal{L}_{\mathrm{LM}}^{\mathrm{CAP}}=\mathbb{E}_{([I;T;O])\sim D}\mathcal{L}_{\mathrm{LM}}([I;T;O])\@add@centering$
@@
In our caption generation stage, we leverage our aligned MLLM to extract entity-related sentences from news articles using multimodal input, without incurring additional model or training costs. Subsequently, our aligned MLLM integrates this information to supplement the original textual input context and generate the news image caption.
Dataset MLLM Setting BLEU-4 ROUGE CIDEr Named entities
P R
GoodNews BLIP-2+GPT-3 Zero_Shot 2.06 13.22 17.12 / /
LLaVA-v1.5 Zero_Shot 2.63 14.03 15.85 17.19 10.96
OSFT 8.48 25.22 78.03 25.69 24.41
MiniGPT-v2 Zero_Shot 1.81 11.81 9.77 15.19 10.17
OSFT 8.05 22.78 62.42 21.32 23.62
InstructBLIP Zero_Shot 2.52 10.78 14.82 13.34 8.78
OSFT 9.53 25.61 78.03 25.89 27.33
NYTimes800k LLaVA-v1.5 Zero_Shot 2.65 13.53 14.66 17.24 11.82
OSFT 7.94 23.48 65.28 25.92 25.01
MiniGPT-v2 Zero_Shot 1.85 11.56 10.21 14.15 11.41
OSFT 7.38 21.27 53.30 21.25 23.89
InstructBLIP Zero_Shot 2.52 11.01 12.74 12.79 11.21
OSFT 10.05 25.45 75.95 27.39 30.37
Table 1. Performance of MLLMs in Zero_Shot and Official supported Supervised Finetune (OSFT) setting. Following the OSFT settings, we train the V-L Connector in MiniGPT-v2 and InstructBLIP, and fully train the V-L Connector together with LLM in LLaVA-v1.5. The result of BLIP-2+GPT-3 is cited from Rajakumar Kalarani et al. (2023). LLaVA-v1.5, MiniGPT-v2 and InstructBLIP are implemented with 7B checkpoints.
3.2. Entity-Aware Multimodal Alignment
@@
Entity-Aware Sentence Selection Previous research frequently entails selecting sentences related to the caption from the provided news article (Zhou et al., 2022a). A common approach involves employing an image-text matching model, such as CLIP, to extract relevant sentences based on news images. However, this method incurs additional training costs for CLIP because it is trained on open-domain datasets, which differ from captions enriched with entities.
@@
Interestingly, the image-text matching task is commonly employed to align visual and textual features. Therefore, we aim to establish a specialized image-text matching task within the news image captioning framework to align vision features with entity-aware textual features. Moreover, leveraging the capability of MLLMs to follow instructions and handle multiple tasks, the process of selecting sentences related to the caption can be seamlessly integrated. This integration not only streamlines the workflow but also reduces the need for separately training a sentence selection model.
@@
Previous research (Zhou et al., 2022a) underscores the significance of entity information in news image captioning tasks. As a result, conventional image-text matching tasks may not be well-suited for entity-aware tasks like news image captioning. Hence, we introduce an Entity-Aware Sentence Selection Task as one of our Entity-Aware Multimodal Alignment tasks.
@@
To construct Entity-Aware Sentence Selection data, we leverage sentences from the news article as the text source. For a given instance comprising the news image $I$, the news article $T={s_{1},s_{2},...,s_{m}}$ consisting of $m$ sentences, and the news image caption $C$, we first identify the visual entities $E=(e_{1},e_{2},...,e_{n})$ mentioned in the caption. Notably, the visual entities exclude entities invisible in the news image, such as ‘DATE’.
@@
(3) $O=\left\{\begin{aligned} &yes,\ \exists e_{j}\in E,s.t.e_{j}\ \in s_{i}\\ &no,\ otherwise\end{aligned}\right.$
For the Entity-Aware Sentence Selection task, each group consists of three sentences containing the most visual entities mentioned in the caption, along with the target caption, serving as positive samples. Additionally, four sentences are chosen from those devoid of any visual entity mentioned in the caption, serving as negative samples.
@@
(4) $\mathcal{L}^{\mathrm{SENT}}_{\mathrm{LM}}=\mathbb{E}_{([I;s_{i};O])\sim D}\mathcal{L}_{\mathrm{LM}}([I;s_{i};O])$
Unlike traditional image-text matching methods, which often randomly select captions corresponding to other images as negative samples, our approach extracts sentences from the same article that do not contain the visual entities as hard negative samples. These sentences frequently present a more akin distribution and cover the same topic, only lacking key entity information. We posit that this strategy affords the model a deeper comprehension of entities.
      Method       BLEU-4       ROUGE       CIDEr       Named entities
      P       R
       GoodNews       Avg+CtxIns(Biten et al., 2019)       0.89       12.20       13.10       8.23       6.06
      ICECAP(Hu et al., 2020)       1.96       15.70       26.08       /       /
      Tell(Tran et al., 2020)       6.05       21.40       53.80       22.20       18.70
      Visual News Captioner(Liu et al., 2021)       6.10       21.60       55.40       22.90       19.30
      JoGANIC(Yang et al., 2021)       6.83       23.05       61.22       26.87       22.05
      NewsMEP(Zhang et al., 2022)       8.30       23.17       63.99       23.43       23.24
      Rajakumar Kalarani et al. (2023)       7.14       24.30       72.33       24.37       20.09
      Qu et al. (2023)       7.91       22.61       69.11       24.38       23.56
      InstructBLIP (OSFT)       9.53       25.61       78.03       25.89       27.33
      EAMA       10.24       27.11       88.39       27.29       28.91
       NYTimes800k       Tell(Tran et al., 2020)       6.30       21.70       54.40       24.60       22.20
      Visual News Captioner(Liu et al., 2021)       6.40       21.90       56.10       24.80       22.30
      JoGANIC(Yang et al., 2021)       6.79       22.80       59.42       28.63       24.49
      NewsMEP(Zhang et al., 2022)       9.57       23.62       65.85       26.61       28.57
      Rajakumar Kalarani et al. (2023)       7.54       23.28       66.41       28.11       23.25
      Qu et al. (2023)       8.81       23.02       70.83       27.57       27.16
      InstructBLIP (OSFT)       10.05       25.45       75.95       27.39       30.37
      EAMA       10.94       27.00       85.61       29.32       31.85
Table 2. Experimental results on GoodNews and NYTimes800k datasets compared with other models. P and R denote Precision and Recall of generated Named entities. InstructBLIP(OSFT) refers to InstructBLIP trained with official scripts on 7B checkpoints. We directly cite the results of models which are not MLLMs from researches by Rajakumar Kalarani et al. (2023); Qu et al. (2023)
@@
(6) $\mathcal{L}^{\mathrm{ENT}}_{\mathrm{LM}}=\mathbb{E}_{([I;T^{*};O_{e}])\sim D}\mathcal{L}_{\mathrm{LM}}([I;T^{*};O_{e}])$
@@
Balance Training Strategy Our alignment training tasks encompass the Entity-Aware Sentence Selection task, the Entity Selection task, and the News Image Captioning task. We concurrently train the model on all three tasks. In line with BLIP2 (Li et al., 2023a), we structure the data into mini-groups, with each group containing a specific number of instances from the News Image Captioning, Entity-Aware Sentence Selection, and Entity Selection tasks.
This ensures that all alignment tasks are trained in every batch, thereby balancing the data flow across all alignment tasks and the loss for each task during the training stage.
All three tasks use the multimodal language modeling loss defined in Equation2, and the final loss is formulated as
(7) $\mathcal{L}=\alpha\mathcal{L}^{\mathrm{SENT}}_{\mathrm{LM}}+\beta\mathcal{L}^{\mathrm{ENT}}_{\mathrm{LM}}+\gamma\mathcal{L}^{\mathrm{CAP}}_{\mathrm{LM}}$
The $\mathcal{L}^{\mathrm{SENT}}_{\mathrm{LM}}$, $\mathcal{L}^{\mathrm{ENT}}_{\mathrm{LM}}$ and $\mathcal{L}^{\mathrm{CAP}}_{\mathrm{LM}}$ are the loss of Entity-Aware Sentence Selection task, Entity Selection task and News Image Captioning task, respectively. Through our Balance Training strategy, our alignment training will primarily enhance the MLLM’s ability on handling multimodal entity information and let the MLLM generate better captions.
3.3. Self-Supplemented Generation

### v5

@@ Abstract
News image captioning requires model to generate an informative caption rich in entities, with the news image and the associated news article. Current MLLMs still bear limitations in handling entity information in news image captioning tasks. Besides, generating high-quality news image captions requires a trade-off between sufficiency and conciseness of textual input information. To explore the potential of MLLMs and address problems we discovered, we propose EAMA: an Entity-Aware Multimodal Alignment based approach for News Image Captioning. Our approach first aligns the MLLM with two extra alignment tasks: Entity-Aware Sentence Selection task and Entity Selection task, together with News Image Captioning task. The aligned MLLM will utilize the additional entity-related information extracted by itself to supplement the textual input while generating news image captions. Our approach achieves better results than all previous models on two mainstream news image captioning datasets.
EAMA : Entity-Aware Multimodal Alignment Based
Approach for News Image Captioning
Junzhe Zhang   Huixuan Zhang   Xunjian Yin   Xiaojun Wan Wangxuan Institute of Computer Technology, Peking University School of Electronics Engineering and Computer Science, Peking University {junzhezhang, zhanghuixuan}@stu.pku.edu.cn {xjyin, wanxiaojun}@pku.edu.cn
1 Introduction
@@
News image captioning, a variant of image captioning task that aims to generate the image caption which contains more specific information, has developed a lot as a standard multimodal task in recent years(Zhang et al., 2022; Rajakumar Kalarani et al., 2023; Qu et al., 2024). In this task, models are required to generate an informative caption, rich in entities such as the names of people or news events, based on the provided news image and associated news article as illustrated in the example shown in Figure 1. According to (Tran et al., 2020), over 96% news image captions contain entity information and approximately 26% words of news image captions are entities in common news image captioning datasets. Compared with general image captioning task, news image captioning task requires model to not only understand the overall information of images, but also apply the knowledge of model to extract related specific information from associated news articles.
@@
Figure 1: An example of news image caption task and responses of the MLLM in different settings. Origin denotes Zero_Shot, OSFT denotes finetuning the MLLM in official settings. EAMA denotes our approach.
@@
Recently, multimodal large language models are developing rapidly and have shown promising capability to address traditional image captioning task. Therefore as a variant of image captioning task, it is natural to consider MLLMs as a promising technique for the news image captioning task. What’s more, news image captioning task requires extracting related information from a long news article, which is exactly a task that can be handled by Large Language models because they contains extensive knowledge and exhibits powerful text processing ability. Since MLLMs mostly contain a LLM component, we believe MLLMs is potentially good at news image captioning task.
@@
However, we find that current MLLMs such as InstructBLIP still face some challenges on news image captioning task: (1) Current MLLMs are not good at dealing with multimodal entity information under zero-shot setting. The ability of generating entities remains insufficient after simply finetuned on news image captioning dataset. An example is shown in Figure 1. (2) While MLLMs are capable of processing long textual information and can take almost entire news articles as textual input, the generation of news image captions still requires a trade-off between sufficiency and conciseness of textual input information.
@@
To address the problems we discovered, we propose EAMA, an Entity-Aware Multimodal Alignment based approach for News Image Captioning. Specifically, we first design two entity-aware alignment tasks, which are Entity-Aware Sentence Selection task and Entity Selection task. Then we align the MLLM on original news image captioning task together with two alignment tasks. Finally, we utilize the aligned MLLM to extract some relevant information from the news article to supplement the textual input context. Our method not only obtains the best result on all automatic metrics but also achieves the competitive performance on entity generation.
@@
Our contributions are  [^1] Our code and data will be released to the community to facilitate future research.:
  • We find that current MLLMs’ ability of dealing with multimodal entity information is insufficient under zero-shot setting and even after simply finetuned. Besides, generating high-quality news image captions with the MLLM still requires a trade-off between sufficiency and conciseness of textual input information.
  • We design two alignment tasks to promote the MLLM’s ability for handling multimodal entity information. The aligned MLLM itself will explicitly extract the entity-related information to supplement textual input context while generating news image captions.
  • Our research conducts comprehensive experiments on two news captioning datasets, GoodNews and NYTimes800k, to demonstrate our analysis. Our proposed method EAMA outperforms all previous methods on both GoodNews and NYTimes800k Datasets.
2 Related Works
2.1 News Image Captioning
News image captioning(Biten et al., 2019; Tran et al., 2020; Liu et al., 2021) is a variant of the general image captioning task which takes news images and associated news articles as input to generate informative captions. In early stage, models(Lu et al. (2018), Avg+CtxIns(Biten et al., 2019)) often generate templates, and then extract entities from news articles. Then, end-to-end methods(Tran et al., 2020; Yang et al., 2021; Zhang et al., 2022; Qu et al., 2024) generate the final caption in one step and achieve higher performance. Rajakumar Kalarani et al. (2023) pretrains OFA model with extra data, and then finetunes OFA on news image captioning datasets. Xu et al. (2024a) adapts finetuned CLIP to construct confidence-aware prompts to fine-tune LLaVA, and construct confidence-feedback prompts to further enhance its capabilities. Xu et al. (2024b) proposes a rule-driven news captioning method, which can generate image descriptions following designated rule signal. As for alignment, Qu et al. (2024) leverages a supervised face-name alignment method to refine face representations, but it only focuses on people’s names. And its alignment method requires modifying model architecture and necessitates that the image contains a facial region. Our approach aligns the MLLM with newly designed alignment tasks focusing on general types of entities with data from the news image caption dataset without additional data.
Since news articles are long for previous models, many methods have been proposed. JoGANIC(Yang et al., 2021) proposes MSTR to read more information. ICECAP(Hu et al., 2020) trains an VSE++ model to retrieve relevant sentences where entities are replaced with categories and then generate the template caption and predict the position of entity tokens in relevant sentences. Zhou et al. (2022a) utilizes finetuned CLIP model and OpenNRE to obtain the key local context. However, MSTR requires much longer input and most retrieval-based methods require an extra cross-modal retrieval module and an extra training cost while the caption model itself remained unimproved. Xu et al. (2024a) also adapts CLIP as an image-text matching module. They construct confidence-aware prompts with CLIP to improve caption model LLaVA. However, they construct potential-positive training samples of CLIP by calculating the textual similarity between the input sentence and the caption. According to previous research Zhou et al. (2022a), captions often contain many non-visual entities, training the image-text matching module CLIP on these data may not be rigorous. The confidence-aware prompts generated by the trained CLIP are also not rigorous to train LLaVA. Our approach aligns the MLLM to promote its ability on handling multimodal entity information. The aligned MLLM can be used to explicitly extract entity-related information to supplement the textual input without extra modules. Our alignment tasks can help model focusing on both visually aware and non-visually aware entity information.
2.2 MLLMs and Alignment
@@
Multimodal large language models(MLLMs) have developed rapidly in recent years. BLIP-2(Li et al., 2023b) applys Q-Former architecture to transform image input into LLMs input tokens. LLaVA(Liu et al., 2023b) and LLaVA-v1.5(Liu et al., 2023a) utilize linear layers or perceptrons to map the vision features into the inputs of LLMs. Through instruction tuning on BLIP2, InstructBLIP(Dai et al., 2023) gains the ability to follow the instruction on different tasks. There are many other MLLMs such as mPLUG-Owl(Ye et al., 2023), Otter(Li et al., 2023a) and Qwen-VL (Bai et al., 2023). Among all MLLMs, GPT-4V(OpenAI et al., 2023) is the most powerful one now. We select some of these MLLMs on our research.
@@
There are numerous studies on multimodal alignment tasks recent years. The most traditional task is image-text matching in CLIP(Radford et al., 2021). There are also fine-grained alignment tasks such as Region-Tag and Region-Noun alignment tasks in Zhou et al. (2022b). MLLMs such as (Li et al., 2023b; Ye et al., 2023) often align textual features and visual features on multimodal tasks with language modeling loss.
3 Approach
3.1 Overview
@@
MLLMs (Multimodal Large Language Models) have been proved highly effective across diverse multimodal tasks, so we craft our methodology based on MLLMs. Figure 2 offers a concise overview, delineating the alignment training and caption generation stages. Typically, MLLM architectures comprise an image encoder, a Vision-Language (V-L) Connector, and a large language model.
@@
During our alignment training stage, we commence by simultaneously training the MLLM on three tasks: Entity-Aware Sentence Selection, Entity Selection, and News Image Captioning. This alignment aims to enhance the MLLM’s capability to handle multimodal entity information.
@@
We employ a language modeling loss as the optimization objective across all tasks. Typically, within a given dataset $D$, each data instance comprises three components: the input image $I$, the textual input context $T$, and the target output text $O=(o_{1},...,o_{n})$, with $o_{i}$ denoting a token. Thus, the general multimodal language modeling loss can be formulated as:
$\mathcal{L}_{\mathrm{LM}}([I;T;O])=-\dfrac{1}{n}\sum\limits_{i=1}^{n}\log\mathbf{P}(o_{i}|o_{1},...,o_{i-1};I;T)$
@@
For example, in News Image Captioning task, the input image $I$ is the news image, the textual input context $T$ is the news article, while the target output text $O$ is the given caption. Thus for a News Image Captioning dataset $D$, the corresponding News Image Captioning loss can be denoted as:
$\centering\mathcal{L}_{\mathrm{LM}}^{\mathrm{CAP}}=\mathbb{E}_{([I;T;O])\sim D}\mathcal{L}_{\mathrm{LM}}([I;T;O])\@add@centering$ (2)
Figure 2: An overview of our approach: EAMA. The left part represents Entity-Aware Multimodal Alignment with Entity-Aware Sentence Selection, Entity Selection and News Image Captioning task. The right part represents the Self-Supplemented Generation for the news image caption given the news article and the news image. The aligned MLLM first extracts related sentences and entities and then generates the news image caption with extracted information as a supplement.
@@
In our caption generation stage, we leverage our aligned MLLM to extract key entities and entity-related sentences from news articles using multimodal input, without incurring additional modules. Subsequently, our aligned MLLM integrates this information to supplement the original textual input context and generate the news image caption.
3.2 Entity-Aware Multimodal Alignment
@@
Entity-Aware Sentence Selection Previous research frequently entails selecting sentences related to the caption from the provided news article (Zhou et al., 2022a). A common approach involves employing an image-text matching model, such as CLIP, to extract relevant sentences based on news images. However, this method incurs additional training stages because CLIP is trained on open-domain datasets, which differ from captions enriched with entities on news image captioning datasets.
@@
Interestingly, the image-text matching task is also commonly employed to align visual and textual features. Therefore, we aim to construct a specialized image-text matching task from news image captioning data to align vision features with entity-aware textual features. Moreover, considering the capability of MLLMs to follow instructions and handle multiple tasks, the process of selecting sentences related to the caption can be seamlessly integrated. Compared with previous methods, this integration not only streamlines the workflow but also reduces the need for separately training an image-text matching module.
@@
Previous research (Zhou et al., 2022a) underscores the significance of entity information in news image captioning tasks. As a result, general image-text matching tasks may not be well-suited for entity-aware tasks like news image captioning. Hence, we introduce an Entity-Aware Sentence Selection task as one of our Entity-Aware Multimodal Alignment tasks.
@@
To construct Entity-Aware Sentence Selection data, we leverage sentences from the news article as the text source. For a given instance comprising the news image $I$, the news article $T={s_{1},s_{2},...,s_{m}}$ consisting of $m$ sentences, and the news image caption $C$, we first identify the visual entities $E=(e_{1},e_{2},...,e_{n})$ mentioned in the caption. Notably, the visual entities exclude some entities invisible in the news image, such as ‘DATE’.
@@
$O=\left\{\begin{aligned} &yes,\ \exists e_{j}\in E,s.t.e_{j}\ \in s_{i}\\ &no,\ otherwise\end{aligned}\right.$ (3)
For the Entity-Aware Sentence Selection task, each group consists of sentences containing the most visual entities mentioned in the caption, along with the target caption, serving as positive samples. Additionally, sentences are chosen from those devoid of any visual entity mentioned in the caption, serving as negative samples.
@@
$\mathcal{L}^{\mathrm{SENT}}_{\mathrm{LM}}=\mathbb{E}_{([I;s_{i};O])\sim D}\mathcal{L}_{\mathrm{LM}}([I;s_{i};O])$ (4)
Unlike traditional image-text matching methods, which often randomly select captions corresponding to other images as negative samples, our approach extracts sentences from the same article that do not contain the visual entities as hard negative samples. These sentences frequently present a more similar distribution and cover the same topic, only lacking visually related entity information. Model can align the multimodal entity information during the alignment task.
@@
$\mathcal{L}^{\mathrm{ENT}}_{\mathrm{LM}}=\mathbb{E}_{([I;T^{*};O_{e}])\sim D}\mathcal{L}_{\mathrm{LM}}([I;T^{*};O_{e}])$ (6)
@@
Figure 3: Performance of MLLMs in Zero_Shot and Official supported Supervised Finetune (OSFT) setting. Following the OSFT settings, we train the V-L Connector in MiniGPT-v2 and InstructBLIP, and fully train the V-L Connector together with LLM in LLaVA-v1.5. LLaVA-v1.5, MiniGPT-v2 and InstructBLIP are implemented with 7B checkpoints.
3.3 Self-Supplemented Generation